1.The volume of a cone is 256 pi in cube and the height is 12 inches. Find the radius
2. Find the diameter of a sphere that has a surface area of 169 pi in squared
3 Find the volume, to the nearest cubic foot, of a sphere whose surface area is 100 ft squared

Answer by mananth(16946)

Volume of cone = 1/3 pi r^2 h
v= 256pi h = 12
256pi = 1/3 pi r^2 *12
3*256pi / pi* 12 = r^2
r^2 = 64
r= 4 inches
..
,..
surface area of sphere = 4pir^2
V= 169 pi
169 pi / 4 pi = r^2
169 /4 = r^2
r= sqrt ( 169/4)
r= 6.5 inches
So diameter = 13 inches
..
..
surface area of sphere = 4pir^2
100/4*pi= r^2
r= 5/sqrt pi
r=2.82
Volume of sphere = 4/3 * pi *r^3
V= (4/3)* pi* (2.82)^3
V= 94 cubic feet